-3x + y + 11z = -410
-4x - 6y + 5z = -214
-2x - 2y + 6z = -244
I need to solve this using any method, but id like to see the steps and the correct answer. I'm not sure where to start with this one, im just lost. Thank you so much for all your help.

Answer by scott8148(6628)

dividing 3rd eqn by -2 gives NEW 3rd eqn __ x+y-3z=122

new 3rd eqn times 3, added to 1st eqn __ 4y+2z=-44 __ 2y+z=-22

new 3rd eqn times 4, added to 2nd eqn __ -2y-7z=274

adding eqns __ -6z=252 __ z=-42

substituting __ 2y+(-42)=-22 __ y=10

substituting __ x+(10)-3(-42)=122 __ x=-14